OpenStudy (anonymous):

The length of an instant message conversation is normally distributed with a mean of 5 minutes and a standard deviation of .7 minutes. What is the probability that a conversation lasts longer than 6 minutes?

OpenStudy (anonymous):

6 minutes is \(\frac{6-5}{.7}=\frac{1}{.7}\) as a "z" score you have to look up in a table to see what percent is above that z score

OpenStudy (anonymous):

\(\frac{1}{.7}=1.4289\) rounded. do you have a table to look this up in?

OpenStudy (anonymous):

yes I do, thank you!

OpenStudy (anonymous):

once you look it up, subtract from 1, because this is the percent that is BELOW that z score, not above it

OpenStudy (anonymous):

who needs a table? use this http://homepage.stat.uiowa.edu/~mbognar/applets/normal.html

OpenStudy (anonymous):

plug in \(\mu =5, x=6, \sigma = .7\) and choose \(P(X>x)\) from the drop down

OpenStudy (iluvsoccer):

@anonymous already did it... :/

OpenStudy (iluvsoccer):

@Elsa213